# Centro de informática, Universidade Federal de Pernambuco, Brazil Hospital das Clínicas, Universidade Federal de Pernambuco. Brazil

### v1

@@ Abstract
AI-based dermatology adoption remains limited by biased datasets, variable image quality, and limited validation. We introduce DermAI, a lightweight, smartphone-based application that enables real-time capture, annotation, and classification of skin lesions during routine consultations. Unlike prior dermoscopy-focused tools, DermAI performs on-device quality checks, and local model adaptation. The DermAI clinical dataset, encompasses a wide range of skin tones, ethinicity and source devices. In preliminary experiments, models trained on public datasets failed to generalize to our samples, while fine-tuning with local data improved performance. These results highlight the importance of standardized, diverse data collection aligned with healthcare needs and oriented to machine learning development.
@@ 1 Introduction
To address these limitations, we introduce a mobile application, DermAI, designed for standardized dermatological image acquisition oriented to AI training and metadata collection in clinical environments. The platform enforces consistent capture guidelines and performs real-time quality checks. Designed to support dermatologists and general practitioners, typically the first point of patient contact, it facilitates early triage and risk assessment while generating clinically meaningful training data. Using the platform, we constructed a new clinical dataset collected during routine consultations at a Brazilian public university hospital, encompassing a wide range of skin tones, ethnicities, devices, and environmental conditions. We benchmarked convolutional neural networks on this dataset and compared their performance to models trained on public datasets; Models trained on public data showed limited generalization under real clinical variability. This result emphasize the need for context-specific samples and demonstrate DermAI as a practical acquisition tool for equitable dermatological AI.
@@ 3.3 Preliminar classification and medical feedback
In the final step, a classification model provides a preliminary suggestion. The model is based on CNNs, following prior work [Pacheco2020-mn, Daneshjou2022-cd]. Several backbones were evaluated, and the best-performing ones were trained and converted to TFLite for local on-device inference. The model classifies $n_{c}$ lesion categories to support clinical decision-making and gather feedback from dermatologists. For general practitioners or non-specialist users, a simplified output is also provided, indicating a binary risk assessment (benign vs. potentially malignant). The classification architecture consists of the CNN backbone followed by average pooling and a fully connected layer trained with categorical cross-entropy. Predictions are shown to clinicians, who can confirm, disagree, or mark uncertainty, entering their own diagnostic hypothesis. This feedback enables performance monitoring under real clinical use. All lesion annotations and metadata remain stored, allowing later updates with biopsy confirmations; lesions flagged as malignant-suspect are prioritized for biopsy follow-up. A supervisory interface allows board-certified dermatologists to review annotation consistency. As the dataset grows, models can be retrained, enabling continuous improvement.
@@ 4 Experiments and discussions
4.1 Assessing classifiers on public datasets
We trained multiple CNN architectures (ConvNeXt, EfficientNet, ResNet, MobileNet, and DenseNet) on the PAD-UFES-20 (PAD) dataset [Pacheco2020-mn], replicating the original setup while adding recent backbones. Since PAD does not define official splits, we established fixed training and test partitions. We then evaluated the resulting models on DDI [Daneshjou2022-cd] and DermAI to assess out-of-distribution generalization. We focused on lightweight models suitable for on-device execution. The DDI, due to its small size, was used mainly for testing. Evaluation was restricted to the six lesion classes provided in PAD (melanoma, nevus, seborrheic keratosis, actinic keratosis, basal cell carcinoma, and squamous cell carcinoma), which permits direct analysis between the three datasets. Performance was measured using accuracy (ACC), recall (R), precision (P), and F1-score (F1) [Pacheco2020-mn]. As shown in Table 1, PAD-trained models align with previous results but show limited generalization when tested on DDI and DermAI, particularly in recall, which is critical for minimizing missed malignant cases. Although PAD is visually diverse, it contains acquisition artifacts (e.g., ink markings, blur) and limited lesion variety or skin tones, which constrains robust feature learning. Retraining the same architectures using DermAI samples improved cross-dataset performance, reflecting the benefits of controlled and diverse acquisition. While it does not reach the highest scores in every scenario, the model maintains stable performance on PAD and preserves useful knowledge when evaluated on DDI, the most challenging dataset given its high noise and variability. We further applied DermAI-based quality filtering to PAD (PAD_filter) and retrained the models. Despite using fewer samples, the PAD_filter improved performance across backbones, highlighting the importance of curated image features. Combining PAD_filter with DermAI yielded the best overall results. Adding DDI samples to training did not improve performance, likely due to their scarcity and noise, reinforcing that increasing data quantity alone is insufficient; quality and curated acquisition is essential for training AI models.
@@ 6 Compliance with ethical standards
This research study was conducted retrospectively using human subject data made available in open access [Pacheco2020-mn, Daneshjou2022-cd]. Ethical approval was not required as confirmed by the license attached with the open access data.

### v2

# Centro de Informática, Universidade Federal de Pernambuco, Brazil Hospital das Clínicas, Universidade Federal de Pernambuco, Brazil

@@ Abstract
AI-based dermatology adoption remains limited by biased datasets, variable image quality, and limited validation. We introduce DermAI, a lightweight, smartphone-based application that enables real-time capture, annotation, and classification of skin lesions during routine consultations. Unlike prior dermoscopy-focused tools, DermAI performs on-device quality checks, and local model adaptation. The DermAI clinical dataset, encompasses a wide range of skin tones, ethinicity and source devices. In preliminary experiments, models trained on public datasets failed to generalize to our samples, while fine-tuning with local data improved performance. These results highlight the importance of standardized, diverse data collection aligned with healthcare needs and oriented to machine learning development. The dataset will be made available for research purposes upon request to the corresponding authors.
@@ 1 Introduction
To address these limitations, we introduce DermAI, designed for standardized dermatological image acquisition oriented to AI training and metadata collection in clinical environments. The platform enforces consistent capture guidelines and performs real-time quality checks. Designed to support dermatologists and general practitioners, it facilitates early triage and risk assessment while generating clinically meaningful training data. Using the platform, we constructed a new clinical dataset collected during routine consultations at a Brazilian public university hospital, encompassing a wide range of skin tones, ethnicities, devices, and environmental conditions. We benchmarked convolutional neural networks on this dataset and compared their performance to models trained on public datasets; Models trained on public data showed limited generalization under real clinical variability. This result emphasize the need for context-specific samples and demonstrate DermAI as a practical acquisition tool.
@@ 3.3 Preliminar classification and medical feedback
In the final step, a classification model provides a preliminary suggestion. The model is based on CNNs, following prior work [Pacheco2020-mn, Daneshjou2022-cd]. Several backbones were evaluated, and the best-performing ones were trained and converted for local on-device inference. The model classifies $n_{c}$ lesion categories to support clinical decision-making and gather feedback from dermatologists. For general practitioners or non-specialist users, the binary risk assessment (benign vs. potentially malignant) is provided. The classification architecture consists of the CNN backbone followed by average pooling and a fully connected layer trained with categorical cross-entropy. Predictions are shown to clinicians, who can confirm, disagree, or indicate uncertainty, and enter their own diagnostic hypothesis. This feedback enables performance monitoring under real clinical use. All lesion annotations and metadata are retained, allowing later updates with biopsy confirmation; lesions flagged as malignant-suspect are prioritized for biopsy follow-up. A supervisory interface enables board-certified dermatologists review annotation consistency.
@@ 4 Experiments and discussions
We trained multiple CNN architectures (Table 1) on the PAD-UFES-20 (PAD) dataset [Pacheco2020-mn], replicating the original setup while adding recent backbones. Since PAD does not define official splits, we established fixed training and test partitions (80/10/10). We then evaluated the resulting models on DDI [Daneshjou2022-cd] and DermAI to assess out-of-distribution generalization. We focused on lightweight models suitable for on-device execution. The DDI, due to its small size, was used mainly for testing. Evaluation was restricted to the six lesion classes provided in PAD (melanoma, nevus, seborrheic keratosis, actinic keratosis, basal cell carcinoma, and squamous cell carcinoma), which permits direct analysis between the three datasets. Performance was measured using accuracy (ACC), recall (R), precision (P), and F1-score (F1) [Pacheco2020-mn]. As shown in Table 1, PAD-trained models align with previous results but show limited generalization when tested on DDI and DermAI, particularly in recall, which is critical for minimizing missed malignant cases. Although PAD is visually diverse, it contains acquisition artifacts (e.g., ink markings, blur) and limited lesion variety or skin tones, which constrain robust feature learning. Retraining the same architectures using DermAI samples improved cross-dataset performance, reflecting the benefits of controlled and diverse acquisition. While it does not achieve the highest scores in every scenario, the model maintains stable performance on PAD and preserves useful knowledge when evaluated on DDI, the most challenging dataset due to its high noise and variability. We further applied DermAI-based quality filtering to PAD (PAD_filter) and retrained the models. Despite using fewer samples, the PAD_filter improved performance across backbones, highlighting the importance of curated image features. Combining PAD_filter with DermAI yielded the best overall results. Adding DDI samples to training did not improve performance, likely due to their scarcity and noise, reinforcing that increasing data quantity alone is insufficient; quality and curated acquisition are essential for training AI models.
@@ 6 Compliance with ethical standards
The study involving human participants was reviewed and approved by the Research Ethics Commitee of the Federal University of Pernambuco (CEP-UFPE), which operates within the national CEP/CONEP system via the Plataforma Brasil platform (CAAE ID 75991523.3.0000.8807). The research was conducted in accordance with the ethical principles embodied in the Declaration of Helsinki and consistent with the Brazilian National Health Council’s Resolution 466/12 (and all applicable local statutory requirements). Written informed consent to participate in this study was obtained from all participants.